# USP: A Unified Sequence Parallelism Approach for Long Context Generative AI

### v4

@@ 2 Sequence Parallelism Approaches
Compared to Tensor Parallelism (TP) [4] and ZeRO [5], the research on SP for Transformer models has been relatively underdeveloped for a long time. The challenge lies in the characteristic of attention computation, where the sequence dimension serves as a common dimension in the $V$ matrix multiplications, making it difficult to partition the tensors and distribute the computation across multiple nodes after slicing the sequence dimension.
Early attempts [6, 7, 8] at SP were not successful, often leading to redundant memory consumption [6] and inefficient communication pattern [8]. For long input sequences, the best practice is to adopt the Sequence Parallelism of Megatron-LM. This method optimizes the AllReduce operation of TP, reducing the memory cost of activations while maintaining the same communication overhead. As shown in Figure 1, the principle of Megatron-LM Sequence Parallelism is the similar to ZeRO-2 [5]. It replaces the AllReduce operation on replicated tensors (the left figure) into equivalent allgather and reduce-scatter operations on partitioned data (the right figure). Since an AllReduce operation is exactly a combination of Allgather and ReduceScatter, the communication cost remains the same. The size of the input and output tensors is reduced by a factor of 1/$N$ across $N$ computational devices. Because of the sequence dimension in input/output tensors is partitioned, it is named as Sequence Parallelism. However, this form of Sequence Parallelism cannot be used independently without tensor parallelism.
@@
The maturity of the standalone SP technology is marked by the publication of two milestone papers in late of 2023. The DeepSpeed-Ulysses [1], named as SP-Ulysses, and Ring-Attention [2], named as SP-Ring, solving the longstanding memory and communication issues inherent to SP from two distinct perspectives. For both methodologies, each computational device is allocated a distinct segment of the $Q$ (query), $K$ (key), $V$ (value), and $O$ (output) tensors, which are segregated along the dimensions of the sequence. There is no redundancy in their storage across devices, which is a primary distinction from the early SP design [6].
SP-Ring can be viewed as a distributed version of FlashAttention [9]. As shown in the right part of Figure 2, SP-Ring employs a two-level nested loop that orchestrates the communication and computation in a blockwise fashion. When computing for the blocks of the tensor $O$ segment, if the required tensor $K$ and $V$ blocks are not locally available, Peer-to-Peer (P2P) communication is utilized to fetch them from other devices. The communication can be organized in a Ring fashion, where each device simultaneously sends and receives $K,V$ blocks, allowing the communication to overlap with computation.
@@
SP-Ulysses leverages All2All communication for segments of the $Q$, $K$, $V$, and $O$ tensors, as shown in the left part of Figure 2. After the All2All operation, the partitioning of these four tensors shifts from the sequence dimension $L$ to the dimension of the attention number heads $hc$. Therefore, the computation of $softmax(QK^{T})V$ for each attention head is maintained in its entirety, and can be implemented using the under the hood Attention operator library, like FlashAttention.
@@ 3 Unified Ulysses-Ring Sequence Parallelism
SP-Ulysses is sensitive to the number of attention heads. The parallelism degree  [^1] The parallelism degree is the number of devices that participate in parallel computation. In other words, it is the number of processes in a parallel process group. of DS-Ulysses cannot exceed the number of attention heads $hc$. Consequently, it is not suitable for GQA (Grouped Query Attention) [10] and MQA (Multi-Query Attention) [11] scenarios. For instance, Llama3-8B employs GQA with a KV head number of 8, which means that when using DS-Ulysses SP, the maximum SP degree is 8. However, if GQA is used and KV head number is 1, DS-Ulysses dose not function. In addition, since Tensor Parallelism also requires division across the $hc$ dimension, SP-Ulysses, and TP are in conflict.
@@
SP-Ring is inefficient in computation and communication. Ring-Attention segments the $Q,K,V,O$ tensors into smaller blocks, which can lead to a decrease in computation efficiency of fused operator $Softmax(QK^{T})V$. Even if the communication and computation fully overlap, the total execution time lags behind that of DS-Ulysses. When using a causal mask has not been addressed, the DS-Ring has load unbalancing issue. DS-Ring does not impose any restrictions on $hc$.
@@
SP-Ulysses and SP-Ring are currently considered alternative strategies for SP, with the choice of one precluding the other. This viewpoint was reinforced by the SP-Ring authors in their ICLR open review rebuttal [^2] https://openreview.net/forum?id=WsRHpHH4s0&noteId=HIY0tae4Gz. Currently, Megatron-DeepSpeed utilizes SP-Ulysses, while Megatron-LM opts for SP-Ring for its SP implementation. However, we claim that rather than viewing them as rivals, they can work together as a unified SP approach.
@@
As shown in Algorithm 1, SP-Ring and SP-Ulysses are organized in a hybrid parallel manner to work together in partitioning the sequence dimension. The SP process group is segmented into two orthogonal process group sets: a set of SP-Ring process groups and a set of SP-Ulysses process groups. For a more intuitive understanding, an SP process group can be viewed as a 2D mesh, SP-Ring operates across each column of the mesh, while SP-Ulysses runs across each row. For example, a process group containing 8 processes could viewed as $2\times 4$, where the SP-Ulysses process group of size 2 and the SP-Ring process group of size 4. This is the same as how data parallelism and tensor parallelism process groups are partitioned.
@@
The inputs to the USP-Attention include the segment of the $Q,K,andV$ tensors after being partitioned along the sequence dimension and the output is a tensor $O$ shard. The size of a tensor segment is ($bs$, $L/N$, $hs$, $hd$). Note that when using MAQ, the shape of the $hc$ for the $K$ and $V$ tensor segments differs from that of the Q tensor. During forward propagation, $scatter\_idx$ is set to 1 and $gather\_idx$ is set to 2. The AllToAll4D merges the $L$ dimension and partitions the $hc$ dimension of the $Q$, $K$, and $V$ tensors, transforming them into ($hc/N$, $bs$, $L$, $d$). They also partition the $O$ tensor along the $L$ dimension and merge the $hc$ dimension. During the backward propagation, $scatter\_idx$ is set to 2, and $gather\_idx$ is set to 1.
@@
The vanilla SP-Ring introduces load-unbalancing issues when applying a causal attention mask, as only the lower triangular matrix needs to be computed. As shown in the Figure 3 If the sequence dimension is divided evenly, the computational tasks are not evenly distributed among the devices. As shown in the left side of the figure, on 4 GPUs the computation load of GPU3 is nearly 7 times that of GPU0.
@@ 4 SP in 4D Parallelism
SP, as a newly emerged parallelism method, how to integrate it into the existing hybrid parallelism framework of Data Parallelism (DP), Tensor Parallelism (TP), and Data Parallelism (PP) has not been thoroughly studied. This section will analyze the relationships between SP and DP, TP, and PP, and discuss the best practices for designing 4D parallelism involving sequence dimension parallelism.
As shown in Table 2, we analyze the communication and memory cost of a standard transformer block for different parallelism. The impact of GQA is not reflected in the table, but we will analyze it later. Entries in the Communications Params columns indicate the collective communication operations are conducted on the parameters and gradients of the transformer block. It includes the parameters/gradients of the weight and bias tensors of 4 Linear Layers in the Self-Attention layer, as well as 2 Linear Layers of the FFN layer, amounting to a volume of $12\times O(d^{2})$ elements in the GPT-2 model. Note that llama3 and llama2 has $12\times O(d^{2})$ elements since the intermediate size is lower. Entries in Communications Act columns indicate the collective communications are conducted on a single hidden states tensor belonging to activations, containing $bs\times L\times d$ elements. The memory cost is broken down into model parameters and gradients (P/G), Optimizer States (OS), and intermediate activation tensors (Act).
@@
SP-Ulysses allreduce $12O(d^{2})$ 8*all2all $8O(bs*L*d)$ $hc$/$L$ $P+G$ $6P$ $A/N$
SP-Ring allreduce $12O(d^{2})$ P2P Overlap $L$/$L$ $P+G$ $6P$ $A/N$
@@
In terms of communication performance, SP is inferior to DP. Both SP and DP require the allreduce operation on gradients during backward propagation. The difference in their communication performance lies in the attention module, where SP introduces additional communications overhead for activations. When the Ulysses degree is set to be greater than 1, the communication overhead of SP will be larger than that of DP due to the all2all operations. When using the Ring method, although the additional P2P communication for attention is overlapped, it introduces extra performance issues. The ideal performance is only to reach the performance of DP without communication for attention. In terms of memory performance, both SP and DP are equivalent, as they all can reduce the activation footprint to 1/$N$.
@@
ZeRO [14] is a distributed parameter management method that reduces the storage space requirements of each computing device by sharding the Optimizer States (ZeRO-1), Gradients (ZeRO-2), and Parameters (ZeRO-3) across multiple devices. The memory cost for Optimizer States, Gradients, and Parameters is reduced to 1/$N$ of the original. ZeRO can also operate within an SP process group because partitioning along the batch dimension ($bs$) or the sequence dimension ($L$) is equivalent to ZeRO’s approach. For ZeRO, SP and DP process groups can be viewed as a combined process group of size $N_{s}p\times N_{d}p$.
@@
The Tensor Parallelism (TP) approach, pioneered by Megatron-LM [4], shards the parameters of models across computing devices. In the TP part of activation tensors, not all are partitioned and distributed across multiple computational devices. Consequently, the memory cost for activations, as reflected in the Table 2, is denoted by the $\alpha A$, where 0<$\alpha$<1. Please refer to the Equation(2) of paper [3] for a more precise $\alpha$. TP has been further refined by Megatron-LM Sequence Parallelism [3], which replaces the an allreduce in TP with an allgather and a reducescatter, and therefore reduces the activation memory cost to $A/P$. To distinguish Megatron-LM Sequence Parallelism from Ring and Ulysses Sequence Parallelism, we named it TP-sp here.
As shown in Table 2, in terms of communication cost, TP is lower than SP. TP and SP-Ulysses have similar communication costs for activations, but TP dose not require to communicate the parameters. However, in long input sequence scenarios, the communication cost of Activations tends to dominate. For the GPT model with the hidden size is 4K, if the input length during training is 8K, 32K, or 64K, the communication cost for Activations is 1.3x, 5.3x, and 10.7x that of the parameter communication cost, respectively. Also, the gradients and parameters communication can be overlapped by computations. We can consider that when dealing with long sequences, the communication cost of TP-sp and SP is similar.
GQA/MQA can reduce SP communication costs, while the communication cost of TP-sp remains unchanged. Assuming the GQA group number is $G$, the Ulysses communication cost for $K$ and $V$ is reduced to 1/$G$, and the activation communication cost is reduced to $4O(bs*L*d)+4O(bs*L*d/G)$. Additionally, SP-Ring, which only communicates $K,V$, also leads to a 1/$G$ reduction in P2P bandwidth demand.
Tip 4: We suggest that SP may have an advantage over TP when employing GQA in terms of communication cost, as GQA can reduce the communication cost of SP without affecting TP.
@@ 6 Future Work
As more and more models transition to a Mixture-of-Experts (MoE) architecture, the research on combining SP with expert parallelism (EP) becomes increasingly significant. Since TP has been proven to be better replaced by EP (Expert Parallelism) [19] due to its inferior communication cost, we argue that SP should also be given a lower priority than EP. This paper does not conduct experiments on MoE models, but we plan to test it in the future.

### v5

@@ 2 Sequence Parallelism Approaches
Compared to Tensor Parallelism (TP) [4] and ZeRO [5], the research on SP for Transformer models has been relatively underdeveloped for a long time. The challenge lies in the characteristic of attention computation, where the sequence dimension serves as a common dimension in the matrix multiplication after softmax, making it difficult to partition the tensors and distribute the computation across multiple nodes after slicing the sequence dimension.
Early attempts [6, 7, 8] at SP were not successful, often leading to redundant memory consumption [6] and inefficient communication pattern [8]. For long input sequences, the best practice is to adopt the Sequence Parallelism of Megatron-LM. This method optimizes the AllReduce operation of TP, reducing the memory cost of activations while maintaining the same communication overhead. As shown in Figure 1, the principle of Megatron-LM Sequence Parallelism is the similar to ZeRO-2 [5]. It replaces the AllReduce operation on replicated tensors (the left figure) into equivalent allgather and reduce-scatter operations on partitioned data (the right figure). Since an AllReduce operation is exactly a combination of Allgather and ReduceScatter, the communication cost remains the same. The size of the input and output tensors is reduced by a factor of 1/$N$ across $N$ computational devices. Because of the sequence dimension in input/output tensors is partitioned, it is named as Sequence Parallelism. However, this form of Sequence Parallelism cannot be used independently without tensor parallelism and communication volume remains constant regardless of the degree of parallelism.
@@
The maturity of the standalone SP technology is marked by the publication of two milestone papers in late 2023. The DeepSpeed-Ulysses [1], named as SP-Ulysses, and Ring-Attention [2], named as SP-Ring, solving the longstanding memory and communication issues inherent to SP from two distinct perspectives. For both methodologies, each computational device is allocated a distinct segment of the $Q$ (query), $K$ (key), $V$ (value), and $O$ (output) tensors, which are segregated along the dimensions of the sequence. There is no redundancy in their storage between devices, which is a primary distinction from the early SP design [6].
SP-Ring can be viewed as a distributed version of FlashAttention [9]. As shown in the right part of Figure 2, SP-Ring employs a nested two-level loop that orchestrates the communication and computation in a blockwise fashion. When computing for the blocks of the tensor $O$ segment, if the required tensor $K$ and $V$ blocks are not locally available, Peer-to-Peer (P2P) communication is utilized to fetch them from other devices. Communication can be organized in a Ring fashion, where each device simultaneously sends and receives $K,V$ blocks, allowing communication to overlap computation.
@@
SP-Ulysses leverages All2All communication for segments of the $Q$, $K$, $V$, and $O$ tensors, as shown in the left part of Figure 2. After the All2All operation, the partitioning of these four tensors changes from the sequence dimension $L$ to the dimension of the attention number heads $hc$. Therefore, the computation of $softmax(QK^{T})V$ for each attention head is maintained in its entirety, and can be implemented using the under the hood Attention operator library, like FlashAttention.
@@ 3 Unified Ulysses-Ring Sequence Parallelism
SP-Ulysses is sensitive to the number of attention heads. The parallelism degree  [^1] The parallelism degree is the number of devices that participate in parallel computation. In other words, it is the number of processes in a parallel process group. of DS-Ulysses cannot exceed the number of attention heads $hc$. Consequently, it is not suitable for the GQA (Grouped Query Attention) [10] and MQA (Multi-Query Attention) [11] scenarios. For instance, Llama3-8B employs GQA with a KV head number of 8, which means that when using DS-Ulysses SP, the maximum SP degree is 8. However, if MQA is used and the KV head number is 1, DS-Ulysses will not function. In addition, since Tensor Parallelism also requires division across the $hc$ dimension, SP-Ulysses, and TP are in conflict.
@@
SP-Ring is inefficient in computation and communication. Ring-Attention segments the $Q,K,V,O$ tensors into smaller blocks, which can lead to a decrease in computation efficiency of the fused operator $Softmax(QK^{T})V$. Even if communication and computation fully overlap, the total execution time lags behind that of DS-Ulysses. When using a causal mask has not been addressed, the DS-Ring has a load-unbalancing issue. DS-Ring does not impose any restrictions on $hc$.
@@
SP-Ulysses and SP-Ring are currently considered alternative strategies for SP, with the choice of one precluding the other. This point of view was reinforced by the SP-Ring authors in their ICLR open review rebuttal [^2] https://openreview.net/forum?id=WsRHpHH4s0&noteId=HIY0tae4Gz. Currently, Megatron-DeepSpeed utilizes SP-Ulysses, while Megatron-LM opts for SP-Ring for its SP implementation. However, we claim that, rather than viewing them as rivals, they can work together as a unified SP approach.
@@
As shown in Algorithm 1, SP-Ring and SP-Ulysses are organized in a hybrid parallel manner named as USP-Attention to work together in partitioning the sequence dimension. The SP process group is segmented into two orthogonal process group sets: a set of SP-Ring process groups and a set of SP-Ulysses process groups. For a more intuitive understanding, an SP process group can be viewed as a 2D mesh, SP-Ring operates across each column of the mesh, while SP-Ulysses runs across each row. For example, a process group containing 8 processes could viewed as $2\times 4$, where a SP-Ulysses process group $ulysses\_pg$ of size 2 and a SP-Ring process group $ulysses\_pg$ of size 4. This is the same as how data parallelism and tensor parallelism process groups are partitioned.
@@
The inputs to the USP-Attention include the segment of the $Q,K,andV$ tensors after being partitioned along the sequence dimension and the output is a tensor $O$ shard. The size of a tensor segment is ($bs$, $L/N$, $hs$, $hd$). Note that when using MAQ, the shape of the $hc$ for the $K$ and $V$ tensor segments differs from that of the Q tensor. During forward propagation, $scatter\_idx$ is set to 1, and $gather\_idx$ is set to 2. AllToAll4D merges the dimension $L$ and partitions the dimension $hc$ of the tensors $Q$, $K$, and $V$, transforming them into ($hc/N$, $bs$, $L$, $d$). They also partition the $O$ tensor along the $L$ dimension and merge the $hc$ dimension. During backward propagation, $scatter\_idx$ is set to 2, and $gather\_idx$ is set to 1.
@@
The vanilla SP-Ring introduces load-unbalancing issues when applying causal attention, as only the lower triangular matrix of $QK^{T}$ needs to be computed. As shown in Figure 3, if the sequence dimension is divided evenly, the computational tasks are not evenly distributed among the devices. As shown in the left side of the figure, on 4 GPUs the computation load of GPU3 is nearly 7 times that of GPU0.
@@ 4 SP in 4D Parallelism
SP, as a newly emerged parallelism method, how to integrate it into the existing hybrid parallelism framework of Data Parallelism (DP), Tensor Parallelism (TP), and Pipeline Parallelism (PP) has not been thoroughly studied. This section will analyze the relationships between SP and DP, TP, and PP, and discuss the best practices for designing 4D parallelism involving sequence dimension parallelism.
As shown in Table 2, we analyze the communication and memory cost of a standard transformer block for different parallelism. The impact of GQA is not reflected in the table, but we will analyze it later. Entries in the Communications Params columns indicate the collective communication operations are conducted on the parameters and gradients of the transformer block. It includes the parameters/gradients of the weight and bias tensors of 4 Linear Layers in the Self-Attention layer, as well as 2 Linear Layers of the FFN layer, amounting to a volume of $12\times O(d^{2})$ elements in the GPT-2 model. Note that llama3 and llama2 has $9.37\times O(d^{2})$ elements since the intermediate size is lower. Entries in Communications Act(ivatiom) columns indicate the collective communications are conducted on a single hidden states tensor belonging to activations, containing $bs\times L\times d$ elements. The memory cost is broken down into model parameters/gradients (P/G), Optimizer States (OS), and intermediate activation tensors (Act).
@@
SP-Ulysses allreduce $12O(d^{2})$ 8*all2all $\frac{8}{N}O(bs*L*d)$ $hc$/$L$ $P+G$ $6P$ $A/N$
SP-Ring allreduce $12O(d^{2})$ P2P $4O(bs*L*d)$ $L$/$L$ $P+G$ $6P$ $A/N$
@@
In terms of communication cost, SP is inferior to DP. Both SP and DP require the allreduce operation on gradients during backward propagation. The difference in their communication performance lies in the attention module, where SP introduces additional communications overhead for activations. When the Ulysses degree is set to be greater than 1, the communication overhead of SP will be larger than that of DP due to the all2all operations. When using the Ring method, although the additional P2P communication for attention is overlapped, it introduces extra performance issues. The ideal performance is only to reach the performance of DP without communication for attention. In terms of memory performance, both SP and DP are equivalent, as they all can reduce the activation footprint to 1/$N$.
@@
ZeRO [14] is a distributed parameter management method that reduces the storage space requirements of each computing device by sharding the Optimizer States (ZeRO-1), Gradients (ZeRO-2), and Parameters (ZeRO-3) across multiple devices. The memory cost for Optimizer States, Gradients, and Parameters is reduced to 1/$N$ of the original. ZeRO can also operate within an SP process group because partitioning along the batch dimension ($bs$) or the sequence dimension ($L$) is equivalent to ZeRO’s approach. ZeRO is working on the unified process group of size $N_{sp}\times N_{dp}$, which combines the SP and DP process groups.
@@
The Tensor Parallelism (TP) approach, pioneered by Megatron-LM [4], shards the parameters of models across computing devices. In the TP part of activation tensors, not all are partitioned and distributed across multiple computational devices. Consequently, the memory cost for activations, as reflected in the Table 2, is denoted by the $\alpha A$, where 0<$\alpha$<1. Please refer to the Equation(2) of paper [3] for a more precise $\alpha$. TP has been further refined by Megatron-LM Sequence Parallelism [3], which replaces the an allreduce in TP with an allgather and a reducescatter, and therefore reduces the activation memory cost to $A/P$ at the cost of redo two allgathers for attention and FFN. To distinguish Megatron-LM Sequence Parallelism from Ring and Ulysses Sequence Parallelism, we named it TP-sp here.
As shown in Table 2, in terms of communication cost, TP-sp is higher than SP-Ulysses and SP-Ring. Firstly, the communication volume of TP-sp is greater than that of SP-Ring, and the latter can be overlapped with computation. Additionally, the communication volume of TP-sp does not decrease with an increase in parallelism, whereas SP-Ulysses can achieve this. Therefore, TP-sp is inferior to any form of sequence parallelism in terms of communication. The SP has a lower communication cost for activations, but it requires synchronizing gradients and parameters. However, the parameter communication volume is small compared to the activation communication volume, and it can be overlapped by computation. GQA/MQA can reduce SP communication costs, while the communication cost of TP-sp remains unchanged. Assuming the GQA group number is $G$, the Ulysses and Ring communication cost for $K$ and $V$ is reduced to 1/$G$, and the activation communication cost is reduced to $\frac{4}{N}O(bs*L*d)+\frac{4}{N}O(bs*L*d/G)$ and $\frac{4}{(}bs*L*\frac{d}{G})$.
Tip 4: We suggest that SP has an advantage over TP-sp in terms of communication cost on a large scale. GQA can further reduce the communication cost of SP.
@@ 5.4 Hybrid SP and TP
5.5 Convergence
We compared the convergence differences between USP and DP. Using the same dataset, we tested the loss curves over 10K iterations on 4 GPUs. We found that the curves for USP and DP completely overlapped, which validates that our modifications to RoPE and SP-Ring for load balancing are correct.
Figure 6: Comparison the loss of DP and USP on LLAMA2-7B with global bs as 4.
@@ 6 Future Work
As more and more models transition to a Mixture-of-Experts (MoE) architecture, the research on combining SP with expert parallelism (EP) becomes increasingly significant. The sequence parallelism employed on the attention module is decoupled with the FFN module. Therefore, hybrid sequence parallelism can also be compatible with Mixture of Experts (MoE), as long as the All2All communication between the Attention and FFN modules is carefully designed.